OpenStudy (anonymous):

Using the distributive property to find the product (y – 4)(y2 + 4y + 16) results in a polynomial of the form y3 + 4y2 + ay – 4y2 – ay – 64. What is the value of a in the polynomial?

OpenStudy (unklerhaukus):

distributive property says \[(A+B)(C+D+\cdots)\\=A(C+D+\cdots)+B(C+D+\cdots)\\ =(AC+AD+A\cdots)+(BC+BD+B\cdots)\]

OpenStudy (anonymous):

??? whats the answer

OpenStudy (unklerhaukus):

(y – 4) (y^2 + 4y + 16) = y (y^2 + 4y + 16) - 4 (y^2 + 4y + 16) =